# To Make, or to Take, That Is the Question: Impact of LOB Mechanics on Natural Trading Strategies

### v1

@@ 2 Data Acquisition
Alternative forms of data. We chose to collect data via live trading experiments due to the limitations of alternative methods, of which there are three broad categories:
@@
    Synthetic orders: This involves the hypothetical placement of synthetic orders within observed market data and tracking their outcomes based on subsequent real market events. The synthetic order placement creates a hypothetical LOB state that includes the order; its potential market impact is ignored, an assumption that is particulary problematic if the order size is large. Even with full L3 data, this method cannot guarantee that a fill is realizable in the real world, and with Binance’s L2 data, accurately tracking an order’s queue position—and thus its outcome—is impossible.
@@
    Passive use of existing orders: This involves observing already-placed orders by other market participants and it is plagued by biases inherent in the orders of other market participants, as discussed in Section 1.1.1. Moreover, this method cannot ensure that the analyzed orders originate from the same trader; in a competitive market with many active participants, the lack of trader identifiability makes this highly unlikely.
@@
Market microstructure and fees. Like most crypto exchanges, Binance employs a tiered fee structure where trading volume is the main factor, though not the only one; under its most favorable tier, the taker fee is 1.5 bp and the maker fee is $-0.5$ bp (i.e., a rebate for maker orders). The tick size is 0.1 USD, or approximately 0.03 bp at the time of the experiment, which is negligible compared to fees. The market generally exhibits robust liquidity, with several hundred thousand USD available at the top bid and ask levels, and a spread that is virtually always one tick wide. In the subsequent analysis, we assume the best possible trading fees; naturally, any strategy results analyzed below would only deteriorate with less favorable fees.
@@ 3 The Two Fundamental Attributes for Maker Orders
We note the caveat that, while latency is a crucial aspect of trading affecting both maker and taker orders, [^8] For a detailed examination of its impact on taker order execution see Albers et al. [2024]; similar effects apply to maker orders. to avoid lengthy technical digressions we treat latency as being negligible. This simplifies our arguments; on the whole, a meticulous treatment of latency reinforces our conclusions.
@@
Taker orders always fill immediately, whereas maker orders have a fill probability less than 1 and fill some time after submission (if at all). Immediate price drift is favorable to taker orders. When a taker buy order is executed, it removes liquidity from the top ask price queue, Where this does not consume the whole queue, it instantaneously increases both the order book imbalance and the microprice, while an order that is large enough to clear the entire queue directly causes the spread to widen upwards, with an increase in the mid-price. Both of these result in an immediate favorable change in the first post-fill order book snapshot, hence the positive immediate price drift (see Equation 2).
@@
It is natural to ask whether this initial momentum persists for longer times. A mid-price move might be counteracted by trades in the opposite direction, for example refilling the queue hit by the taker order that causes the spread to widen with associated immediate price drift. However, as widely acknowledged in the literature Lipton et al. [2013], Gould and Bonart [2015], Lehalle and Mounjid [2016], Stoikov [2017], there is a positive correlation between order book imbalance and subsequent asset price movements, typically observed over short time scales ranging from milliseconds to seconds. Moreover, if the taker order causes an immediate mid-price movement, the well-documented autocorrelation of high-frequency returns suggests that this trend is likely to continue in the short term. Our empirical evidence confirms these statements, and in Section  7, we explore why the aforementioned correlations are less than 1: the fill-producing initial momentum occasionally reverts. This class of cases is of paramount importance in market making, as we shall see later.
In summary, maker orders experience negative instantaneous price drift and negative expected price drift over longer time scales, amounting to adverse selection.
@@ 4 Fill Probabilities
Suppose $\mathfrak{o}$ is a maker order submitted at time $t_{0}$ to the near-side queue, and consider how its fill probability depends on the $\operatorname{LA}_{t}(\mathfrak{o})$ on the near side and the opposite-side queue size $Q_{t}^{\text{opp}}$. We present two informal arguments, the first concerning $\operatorname{LA}_{t}(\mathfrak{o})$ and the second concerning $Q_{t}^{\text{opp}}$, and validate their predictions against our empirical data.
1: Near-side liquidity. The order fills at time $T(\mathfrak{o})>t_{0}$ if and only if at that time a sell taker order is processed within the exchange matching engine and its quantity exceeds the order’s liquidity ahead, $\operatorname{LA}_{T}(\mathfrak{o})$, consuming $\mathfrak{o}$ and all maker orders earlier in the queue. Since the arrival probability of a taker order with size less than a given quantity decreases as that quantity decreases, for times $t_{0}<t\leq T$ we would expect less liquidity ahead $\operatorname{LA}_{t}(\mathfrak{o})$ to yield a higher fill probability. Furthermore, since $\operatorname{LA}_{t}(\mathfrak{o})\leq\operatorname{LA}_{t_{0}}(\mathfrak{o})$, the order $\mathfrak{o}$ certainly fills if at any point after its submission (and before a potential cancellation), a taker order exceeding $\operatorname{LA}_{t_{0}}(\mathfrak{o})$ arrives. Therefore, we would generally expect orders posted on a small near-side queue to have a high likelihood of filling.
@@
We now examine how fill frequencies (sometimes also referred to as empirical fill probabilities) depend empirically on the conditions at the order’s submission time $t_{0}(\mathfrak{o})$. Specifically, we analyze the fill frequency conditioned on the initial near-side queue size $Q^{\text{near}}_{t_{0}}$ (equivalent to the order’s initial liquidity ahead) and the initial opposite-side queue size $Q^{\text{opp}}_{t_{0}}$, namely the probability
@@
To further quantify the relationship between fill probability, queue sizes, and imbalance, we fitted an ordinary least-squares (OLS) regression model to the surface shown in Figure 4. The regression model takes the form:
@@
where $z$ represents the fill probability, $Q^{\text{near}}$ and $Q^{\text{opp}}$ denote the opposite-side and near-side queue sizes, and $\text{imb}\vcentcolon=(Q^{\text{near}}-Q^{\text{opp}})/(Q^{\text{near}}+Q^{\text{opp}})$ represents the imbalance. The model achieves an $R^{2}$ of 0.946, indicating a near-perfect fit to the data, and underscoring that fill probabilities can be estimated with minimal complexity.
The estimated coefficients are $\beta_{0}=0.5649$, $\beta_{1}=0.0159$, $\beta_{2}=0.1013$, and $\beta_{3}=-0.3166$. The p-values show that all terms are statistically significant except $Q^{\text{opp}}$ ($P=0.065$), with the dominant drivers being the near-side queue size ($P=0.000$) and imbalance ($P=0.000$).
@@
A basic “zero-intelligence" approach, ignoring $\operatorname{LB}_{t}(\mathfrak{o})$ and potential responses of other traders to the size of the whole near-side queue, including $\operatorname{LB}_{t}(\mathfrak{o})$, might suggest that the fill probability of $\mathfrak{o}$ at time $t$ is the same as that of another order $\mathfrak{o}^{\prime}$ at its initial time $t_{0}(\mathfrak{o^{\prime}})$ if they have the same liquidity ahead: $\left(\operatorname{LA}_{t}(\mathfrak{o}),Q_{t}^{\text{opp}}\right)=\left(\operatorname{LA}_{t_{0}(\mathfrak{o^{\prime}})}(\mathfrak{o^{\prime}}),Q_{t_{0}(\mathfrak{o^{\prime}})}^{\text{opp}}\right).$ The rationale is that the total quantity of liquidity removal necessary to terminate the order, either on the near-side (giving a fill) or on the opposite-side (giving a cancellation) is the same for both orders. However, a build-up of $\operatorname{LB}_{t}(\mathfrak{o})$ is likely to reduce the arrival rate of taker orders on the near-side, increase it on the opposite side, reduce cancellations from the near-side (particularly for front-of-queue orders), and increase the cancellation rate of opposite-side orders, as traders respond to the evolving imbalance. All of these reactions tend to reduce the fill probability for $\mathfrak{o}$ compared with $\mathfrak{o}^{\prime}$. This estimate is thus an upper bound on the order’s fill probability at the intermediate time. The deviation from the true fill probability is determined by the extent to which a non-zero $\operatorname{LB}_{t}(\mathfrak{o})$ influences the decisions of other traders.
@@ 5 Price Drift
2: The opposite-side queue. Now consider the opposite-side queue size $Q^{\text{opp}}_{T}$ at fill time. If $Q^{\text{opp}}_{T}$ is very small, then even a small trade suffices to produce a post-fill price change favorable to the order $\mathfrak{o}$. Conversely, as $Q^{\text{opp}}_{T}$ increases, this post-fill price move becomes increasingly less likely. Altogether, we would therefore expect the degree of adverse selection—measured by the conditinoal expected markout return—to worsen as the opposite-side queue size increases.
@@
Building on this, we define bins for $\operatorname{QP}_{T}$ based on percentile ranges. For instance, 0% - 10% ($\operatorname{QP}_{T}\in[0,0.1]$) represents orders within the front 10% in the queue at the time of execution, while 75% - 100% ($\operatorname{QP}_{T}\in[0.75,1]$) represents orders within the last quartile in the queue. We then examine the empirical markout return distributions conditioned on a near-side queue size bin, an opposite-side queue size bin, and a queue position bin.
@@
The histograms reveal a striking difference in outcomes. For fills with a front-of-queue position, the distribution’s mass is predominantly at positive values, with a notable spike at small positive markout returns, as reflected in the average value of $-0.058$ bp. Conversely, fills from maker orders at the back of the queue predominantly result in negative outcomes, with the majority of the distribution’s mass being at negative values, and an average return of $-0.775bp$. Notably, the standard deviation of markout returns is also significantly larger for back-of-queue fills, indicating greater variability in outcomes. These observations suggest that trading in the direction of the order book imbalance with a maker order is only desirable if the order is filled from a front-of-queue position. [^10] As we shall see later, trading in the direction of the imbalance with a taker order does not generally outperform the taker fee, rendering it a losing strategy overall.
@@
Markout return histograms for orders where both $Q^{\text{near}}_{T}$ and $Q^{\text{opp}}_{T}$ fall into the “large" bin are shown in Figure 8 for fills from orders within the 0%-10% front-of-queue position range, while those within the 75%-100% range are shown in Figure 8. Consistent with the previous pair of histograms, we see significantly better outcomes for front-of-queue orders. The outcomes are slightly worse on average (see Table 1) compared with the previous case, consistent with the less favorable order book imbalance in this scenario: the average markout return is $-0.296$ bp for front-of-queue fills and $-1.157$ bp for back-of-queue fills. The back-of-queue histogram, in particular, shows that the vast majority of outcomes are adverse, with almost all fills resulting in negative markout returns.
@@
The last pair of histograms in the main text (additional histograms are available in the appendix) show markout returns for orders for which $Q^{\text{near}}_{T}$ is small and $Q^{\text{opp}}_{T}$ is large, indicating that the order book imbalance is unfavorable. Outcomes for front-of-queue positions are shown in Figure 10, with an average markout return of $-0.539$ bp as reflected in Table 1, while those for back-of-queue positions are shown in Figure 10, where the average markout return is slightly worse at $-0.763$ bp. Most outcomes in both cases are adverse, with markout returns predominantly negative. Interestingly, the difference between the two histograms is not significant, suggesting that when the queue is small, the queue position does not matter much. This aligns with our previous discussion and makes economic sense: if the queue is small, the liquidity behind an order at the beginning of the queue is not much different from that of an order at the back of the queue.
@@
Large near, small opp. 0 - 0.1 $-0.058$ $-9.2$ $6.3$ $1.0$
0.1 - 0.4 $-0.586$ $-9.5$ $7.0$ $1.3$
0.4 - 0.75 $-0.743$ $-8.6$ $6.0$ $1.3$
0.75 - 1 $-0.775$ $-27.3$ $9.5$ $1.6$
Large near, large opp. 0 - 0.1 $-0.296$ $-8.9$ $4.1$ $0.9$
0.1 - 0.4 $-0.882$ $-6.1$ $3.0$ $1.1$
0.4 - 0.75 $-0.967$ $-7.2$ $3.9$ $1.1$
0.75 - 1 $-1.157$ $-5.8$ $8.0$ $1.1$
Small near, small opp. 0 - 0.1 $-0.562$ $-16.7$ $18.8$ $1.6$
0.1 - 0.4 $-0.711$ $-13.8$ $8.8$ $1.8$
0.4 - 0.75 $-0.622$ $-14.0$ $17.3$ $1.9$
0.75 - 1 $-0.677$ $-19.8$ $20.2$ $2.2$
Small near, large opp. 0 - 0.1 $-0.539$ $-19.0$ $21.0$ $0.8$
0.1 - 0.4 $-0.645$ $-33.4$ $5.2$ $0.9$
0.4 - 0.75 $-0.686$ $-9.8$ $9.7$ $0.8$
0.75 - 1 $-0.763$ $-14.6$ $10.7$ $0.9$

### v2

@@ 2  Data Acquisition
Alternative forms of data. A key contribution of this paper is its experimental methodology. We chose to collect data via live trading experiments to obtain a signal-bias-free dataset, overcoming the inherent limitations of alternative methods, of which there are three broad categories:
@@
    Synthetic orders: This involves the hypothetical placement of synthetic orders within observed market data and tracking their outcomes based on subsequent real market events. The synthetic order placement creates a hypothetical LOB state that includes the order; its potential market impact is ignored, an assumption that is particularly problematic if the order size is large. Even with full L3 data, this method cannot guarantee that a fill is realizable in the real world, and with Binance’s L2 data, accurately tracking an order’s queue position—and thus its outcome—is impossible.
@@
    Passive use of existing orders: This involves observing already-placed orders by other market participants and it is subject to biases inherent in the orders of other market participants, as discussed in Section 1.1.1. Moreover, this method cannot ensure that the analyzed orders originate from the same trader; in a competitive market with many active participants, the lack of trader identifiability makes this highly unlikely.
@@
Market microstructure and fees. Like most crypto exchanges, Binance employs a tiered fee structure where trading volume is the main factor, though not the only one; under its most favorable tier, the taker fee is 1.5 bp and the maker fee is 0.5-0.5 bp (i.e., a rebate for maker orders). The tick size is 0.1 USD, or approximately 0.03 bp at the time of the experiment, which is negligible compared to fees. The market generally exhibits robust liquidity, with several hundred thousand USD available at the top bid and ask levels, and a spread that is virtually always one tick wide. In the subsequent analysis, we assume the best possible trading fees; naturally, any strategy results analyzed below would only deteriorate with less favorable fees.
@@ 3  The Two Fundamental Attributes for Maker Orders
We note the caveat that, while latency is a crucial aspect of trading affecting both maker and taker orders, [^10] For a detailed examination of its impact on taker order execution see Albers_2024; similar effects apply to maker orders. to avoid lengthy technical digressions we treat latency as being negligible. This simplifies our arguments; on the whole, a meticulous treatment of latency reinforces our conclusions.
@@
Taker orders always fill immediately, whereas maker orders have a fill probability less than 1 and fill some time after submission (if at all). Immediate price drift is favorable to taker orders. When a taker buy order is executed, it removes liquidity from the top ask price queue; where this does not consume the whole queue, it instantaneously increases both the order book imbalance and the microprice, while an order that is large enough to clear the entire queue directly causes the spread to widen upwards, with an increase in the mid-price. Both of these result in an immediate favorable change in the first post-fill order book snapshot, hence the positive immediate price drift (see Equation 2).
@@
It is natural to ask whether this initial movement persists for longer times. A mid-price move might be counteracted by trades in the opposite direction, for example refilling the queue hit by the taker order that first caused the spread to widen with associated immediate price drift. However, as widely acknowledged in the literature Lipton:2013aa; Gould:2015aa; Lehalle:2016aa; Stoikov_2017, there is a positive correlation between order book imbalance and subsequent asset price movements, typically observed over short time scales ranging from milliseconds to seconds. Moreover, if the taker order causes an immediate mid-price movement, the well-documented autocorrelation of high-frequency returns suggests that this trend is likely to continue in the short term. Our empirical evidence confirms these statements, and in Section  7, we explore why the correlations involved are less than 1: the fill-producing initial movement occasionally reverts. This class of cases is of paramount importance in market making, as we shall see later. In summary, maker orders experience negative instantaneous price drift and negative expected price drift over longer time scales, amounting to adverse selection.
@@ 4  Fill Probabilities
Suppose $\mathfrak{o}$ is a maker order submitted at time $t_{0}$ to the near-side queue, and consider how its fill probability depends on $\operatorname{LA}_{t}(\mathfrak{o})$ on the near side and the opposite-side queue size $Q_{t}^{\text{opp}}$. We present two informal arguments, the first concerning $\operatorname{LA}_{t}(\mathfrak{o})$ and the second concerning $Q_{t}^{\text{opp}}$, and validate their predictions against our empirical data.
1: Near-side liquidity. The order fills at time $T(\mathfrak{o})>t_{0}$ if and only if at that time a sell taker order is processed within the exchange matching engine and its quantity exceeds the order’s liquidity ahead, $\operatorname{LA}_{T}(\mathfrak{o})$, consuming $\mathfrak{o}$ and all maker orders earlier in the queue. Since the arrival probability of a taker order with size less than a given quantity decreases as that quantity decreases, for times $t_{0}<t\leq T$ we would expect less liquidity ahead to yield a higher fill probability. Furthermore, since $\operatorname{LA}_{t}(\mathfrak{o})\leq\operatorname{LA}_{t_{0}}(\mathfrak{o})$, the order $\mathfrak{o}$ is certain to fill if, while it remains active in the book, a taker order arrives at its price level with size exceeding $\operatorname{LA}_{t_{0}}(\mathfrak{o})$. Therefore, we would generally expect orders posted on a small near-side queue to have a high likelihood of filling.
@@
We now examine how fill frequencies (sometimes also referred to as empirical fill probabilities) depend empirically on the conditions at the order’s submission time $t_{0}(\mathfrak{o})$. Specifically, we analyze the fill frequency conditioned on the near-side queue size at the time of submission $Q^{\text{near}}_{t_{0}}$ (equivalent to the order’s initial liquidity ahead) and the initial opposite-side queue size $Q^{\text{opp}}_{t_{0}}$; namely, the probability
@@
To further quantify the relationship between fill probability, queue sizes, and imbalance, we fitted an ordinary least-squares (OLS) regression model to the surface shown in Figure 4. For this analysis, the queue size variables were normalized (by dividing by their 99th percentile values) to ensure their coefficients are on a comparable scale. The regression model takes the form:
@@
where $z$ represents the fill probability, $Q^{\text{near}}$ and $Q^{\text{opp}}$ denote the normalized opposite-side and near-side queue sizes (with a slight abuse of notation, we reuse these symbols for simplicity), and $\text{imb}\vcentcolon=(Q^{\text{near}}-Q^{\text{opp}})/(Q^{\text{near}}+Q^{\text{opp}})$ represents the imbalance. The model achieves an $R^{2}$ of 0.946, indicating a near-perfect fit to the data, and underscoring that fill probabilities can be estimated with minimal complexity.
The estimated coefficients are $\beta_{0}=0.5649$, $\beta_{1}=0.0159$, $\beta_{2}=0.1013$, and $\beta_{3}=-0.3166$. The p-values show that all terms are statistically significant except $Q^{\text{opp}}$ ($\mathrm{p}=0.065$), with the dominant drivers being the near-side queue size ($\mathrm{p}=0.000$) and imbalance ($\mathrm{p}=0.000$).
@@
A basic “zero-intelligence” approach, ignoring $\operatorname{LB}_{t}(\mathfrak{o})$ and potential responses of other traders to the size of the whole near-side queue, including $\operatorname{LB}_{t}(\mathfrak{o})$, might suggest that the fill probability of $\mathfrak{o}$ at time $t$ is the same as that of another order $\mathfrak{o}^{\prime}$ at its initial time $t_{0}(\mathfrak{o^{\prime}})$ if they have the same liquidity ahead: $\left(\operatorname{LA}_{t}(\mathfrak{o}),Q_{t}^{\text{opp}}\right)=\left(\operatorname{LA}_{t_{0}(\mathfrak{o^{\prime}})}(\mathfrak{o^{\prime}}),Q_{t_{0}(\mathfrak{o^{\prime}})}^{\text{opp}}\right).$ The rationale is that the total quantity of liquidity removal necessary to terminate the order, either on the near-side (giving a fill) or on the opposite-side (giving a cancellation) is the same for both orders. However, a build-up of $\operatorname{LB}_{t}(\mathfrak{o})$ is likely to reduce the arrival rate of taker orders on the near-side, increase it on the opposite side, reduce cancellations from the near-side (particularly for front-of-queue orders), and increase the cancellation rate of opposite-side orders, as traders respond to the evolving imbalance. All of these reactions tend to reduce the fill probability for $\mathfrak{o}$ compared with $\mathfrak{o}^{\prime}$. This estimate is thus an upper bound on the order’s fill probability at the intermediate time. The deviation from the true fill probability is determined by the extent to which a non-zero $\operatorname{LB}_{t}(\mathfrak{o})$ influences the decisions of other traders.
@@ 5  Price Drift
2: The opposite-side queue. Now consider the opposite-side queue size $Q^{\text{opp}}_{T}$ at fill time. If $Q^{\text{opp}}_{T}$ is very small, then even a small trade suffices to produce a post-fill price change favorable to the order $\mathfrak{o}$. Conversely, as $Q^{\text{opp}}_{T}$ increases, this post-fill price move becomes increasingly less likely. Altogether, we would therefore expect the degree of adverse selection—measured by the conditional expected markout return—to worsen as the opposite-side queue size increases.
@@
Building on this, we define bins for $\operatorname{QP}_{T}$ based on percentile ranges. For instance, 0% - 10% ($\operatorname{QP}_{T}\in[0,0.1)$) represents orders within the front 10% in the queue at the time of execution, while 75% - 100% ($\operatorname{QP}_{T}\in[0.75,1]$) represents orders within the last quartile in the queue. We then examine the empirical markout return distributions conditioned on a near-side queue size bin, an opposite-side queue size bin, and a queue position bin.
@@
The histograms reveal a striking difference in outcomes. For fills with a front-of-queue position, the distribution’s mass is predominantly at positive values, with a notable spike at small positive markout returns, as reflected in the average value of 0.058-0.058 bp. Conversely, fills from maker orders at the back of the queue predominantly result in negative outcomes, with the majority of the distribution’s mass being at negative values, and an average return of 0.775-0.775 bp. Notably, the standard deviation of markout returns is also significantly larger for back-of-queue fills, indicating greater variability in outcomes. These observations suggest that trading in the direction of the order book imbalance with a maker order is only desirable if the order is filled from a front-of-queue position. [^12] As we shall see later, trading in the direction of the imbalance with a taker order does not generally outperform the taker fee, rendering it a losing strategy overall.
@@
Markout return histograms for orders where both $Q^{\text{near}}_{T}$ and $Q^{\text{opp}}_{T}$ fall into the “large” bin are shown in Figure 8 for fills from orders within the 0%-10% front-of-queue position range, while those within the 75%-100% range are shown in Figure 8. Consistent with the previous pair of histograms, we see significantly better outcomes for front-of-queue orders. The outcomes are slightly worse on average (see Table 1) compared with the previous case, consistent with the less favorable order book imbalance in this scenario: the average markout return is 0.296-0.296 bp for front-of-queue fills and 1.157-1.157 bp for back-of-queue fills. The back-of-queue histogram, in particular, shows that the vast majority of outcomes are adverse, with almost all fills resulting in negative markout returns.
@@
The last pair of histograms in the main text (additional histograms are available in the appendix) show markout returns for orders for which $Q^{\text{near}}_{T}$ is small and $Q^{\text{opp}}_{T}$ is large, indicating that the order book imbalance is unfavorable. Outcomes for front-of-queue positions are shown in Figure 10, with an average markout return of 0.539-0.539 bp as reflected in Table 1, while those for back-of-queue positions are shown in Figure 10, where the average markout return is slightly worse at 0.763-0.763 bp. Most outcomes in both cases are adverse, with markout returns predominantly negative. Interestingly, the difference between the two histograms is not significant, suggesting that when the queue is small, the queue position does not matter much. This aligns with our previous discussion and makes economic sense: if the queue is small, the liquidity behind an order at the beginning of the queue is not much different from that of an order at the back of the queue.
@@
Large near, small opp. 0 - 0.1 0.058-0.058 9.2-9.2 $6.3$ $1.0$
0.1 - 0.4 0.586-0.586 9.5-9.5 $7.0$ $1.3$
0.4 - 0.75 0.743-0.743 8.6-8.6 $6.0$ $1.3$
0.75 - 1 0.775-0.775 27.3-27.3 $9.5$ $1.6$
Large near, large opp. 0 - 0.1 0.296-0.296 8.9-8.9 $4.1$ $0.9$
0.1 - 0.4 0.882-0.882 6.1-6.1 $3.0$ $1.1$
0.4 - 0.75 0.967-0.967 7.2-7.2 $3.9$ $1.1$
0.75 - 1 1.157-1.157 5.8-5.8 $8.0$ $1.1$
Small near, small opp. 0 - 0.1 0.562-0.562 16.7-16.7 $18.8$ $1.6$
0.1 - 0.4 0.711-0.711 13.8-13.8 $8.8$ $1.8$
0.4 - 0.75 0.622-0.622 14.0-14.0 $17.3$ $1.9$
0.75 - 1 0.677-0.677 19.8-19.8 $20.2$ $2.2$
Small near, large opp. 0 - 0.1 0.539-0.539 19.0-19.0 $21.0$ $0.8$
0.1 - 0.4 0.645-0.645 33.4-33.4 $5.2$ $0.9$
0.4 - 0.75 0.686-0.686 9.8-9.8 $9.7$ $0.8$
0.75 - 1 0.763-0.763 14.6-14.6 $10.7$ $0.9$